If one solution is 20% acid and another 50% acid, then how many liters of each must be mixed to get 20 liters of solution that is 40% acid?
Answer by richwmiller(17219)

We need a total of 20 liter at 40%
We want to know how many liter at 20% to mix with 50%
a+b=20,
0.2*a+0.5*b=0.4*20
a=20-b
0.2*(20-b)+0.5*b=8
4-0.2b+0.5*b=8
0.3*b=4
b=13.3333333
a=20-b
a=6.66666667 liter at 20%
b=13.3333333 liter at 50%
check
0.2*6.66666667+0.5*13.3333333=0.4*20
1.33333333+6.66666667=8
8.0=8
ok
codemx1